OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 49x2 - 28x + 16 4a2 - 20a + 25 25b2 - 20b - 16 16x2 - 24x - 9

OpenStudy (anonymous):

\[(2a-5)^2=4a^2-20a+25\] so that one works as well maybe they all are?

OpenStudy (anonymous):

thank you. can you help me on another one?

OpenStudy (anonymous):

i better be more careful here \[(7x-4)^2=49x^2-28x-28x+16=49x^2-56x+16\] so NO to the first one

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

lets do these first, then we can do another one

OpenStudy (anonymous):

sorry just thought you were done with the first one :D

OpenStudy (anonymous):

the only one here that is a perfect square is \[4a^2 - 20a + 25\] because \[(2a-5)^2=(2a-5)(2a-5)=4a^2-10a-1a+25=4a^2-20a+25\]

OpenStudy (anonymous):

only the second choice is right ok now we can do another

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Some steps to rewrite the expression x3 - 9x + x2 - 9 as a product of three factors are shown below: Step 1: x3 - 9x + x2 - 9 Step 2: x3 + x2 - 9x - 9 Step 3: x2(x + 1) - 9(x + 1) Which of the following best shows the next two steps to rewrite the expression? Step 4: (x2 + 9)(x + 1); Step 5: (x + 3)(x + 3)(x + 1) Step 4: (x2 - 9)(x + 1); Step 5: (x + 3)(x + 3)(x + 1) Step 4: (x2 + 9)(x + 1); Step 5: (x - 3)(x + 3)(x + 1) Step 4: (x2 - 9)(x + 1); Step 5: (x - 3)(x + 3)(x + 1)

OpenStudy (anonymous):

we start here \[ x^2(x + 1) - 9(x + 1)\] the common factor is \((x+1)\) so step 4 is \[(x^2-9)(x+1)\] and then step 5 is \[(x-3)(x+3)(x+1)\] now lets look at your choices

OpenStudy (anonymous):

last one has \[Step 4: (x^2 - 9)(x + 1)\\Step 5: (x - 3)(x + 3)(x + 1)\] pick that one

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw